OpenStudy (anonymous):

x^2+8y=0 find the vertex focus and directrix

OpenStudy (anonymous):

All of these are straight formula vertex (-b/2a, (4ac-b^2)/(4a)

OpenStudy (anonymous):

u losing me

OpenStudy (anonymous):

First you have to solve for y

OpenStudy (anonymous):

This is your go-by:\[y=ax ^{2}+bx=c\]That is how you determine what a and b is in above formula. The formula in the previous post represent (h,k) the vertex. Sorry for all the alphabet soup. Don't get lost, stay with me.

OpenStudy (anonymous):

y didnt u use (x-h)^2=4p(y-k)

OpenStudy (anonymous):

i know the answer to the problem i just dont understand how they got it

OpenStudy (anonymous):

the answer is (0,0) focus is (0,-2) and the directrix is y=2 so how they get that

OpenStudy (anonymous):

That is a valid method. If you don't understand that one. Try the one I posted above. It looks simpler to me.

OpenStudy (anonymous):

how do i do the above one when i have no a and b

OpenStudy (anonymous):

This is how you do it. I'll walk you through it. First solve for y. Put it in the form y=

OpenStudy (anonymous):

i got -x^2/8=y

OpenStudy (anonymous):

OK, from the formula, write your a, b, c

OpenStudy (anonymous):

u lost me how do i do that to me they not relating im real smart but sometimes u gotta talk 2 me like im slow

OpenStudy (anonymous):

I miswrote it the first time\[y=ax ^{2}+bx+c\]Your equation is\[y=(-1/8)x ^{2}+0x+0\]

OpenStudy (anonymous):

ok so thats used 2 find what

OpenStudy (anonymous):

First post message

OpenStudy (anonymous):

ok so is b and c the vertex points

OpenStudy (anonymous):

You choose your a, b, c. a=-1/8, b=0, c=0. Put it in the formula {(-0/2(-1), [4(-1)(0)-0^2]/[4(-1)]} which is (0,0). (h,k) is (0,0)

OpenStudy (anonymous):

ok so the long w u find the vertex and u use the other one to help find wast goes in the 1st formula correct???

OpenStudy (anonymous):

I'm sorry. I didn't get that.

OpenStudy (anonymous):

the formula u wrote at the very 1st post is that the one i need to get the vertex but i formula to get the numbers that go in the first formula

OpenStudy (anonymous):

dat made no sense lol

OpenStudy (anonymous):

ok both formulas u gave me is that how i get the vertext??

OpenStudy (anonymous):

The vertex is two numbers (0,0) So the formula represent the two numbers.

OpenStudy (anonymous):

i dont understand how you did that 1 st formula its 2 much going on wit all th () and {} [] like no how r u doing dat the vertex is (0,0) now explain to me where in the formula i get dat from cuz trying to solve that formula looks impossible

OpenStudy (anonymous):

That is from the first formula, the first message in the post, but if you look at it long enough, you would notice your b=0, c=0, so quickly everything else ends up 0.

OpenStudy (anonymous):

i said dat like 7 post ago ur b and c is ur h and k which is he vertex right??

OpenStudy (anonymous):

Well not exactly. It just so happened that the zeros dominated, you could very well have a=1, b=2, c=3. That doesn't mean h=2, k=3

OpenStudy (anonymous):

uggggggggggggggghhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhh ok how do u get the focus point

OpenStudy (anonymous):

This might help http://www.youtube.com/watch?v=ZJf9shWlMz0

OpenStudy (anonymous):

not helping

OpenStudy (anonymous):

ok can u start fresh with me step by step??

OpenStudy (anonymous):

I'm not up on all that. I just had that vertex info handy in a book I have around the house. Hang out at you Math lab at school.

OpenStudy (anonymous):

not bout 2 happen ill just give up but thanks